OpenStudy (anonymous):

please help... Using one of the laws of exponents, prove that any number raised to the 0'th power equals 1

OpenStudy (zzr0ck3r):

1=X^1/x^1=x^(1-1)=x^0

OpenStudy (zzr0ck3r):

This not true for 0^0 so the question is not right. This could also be used to show 0^0 is undefined

OpenStudy (zzr0ck3r):

Understand?

OpenStudy (anonymous):

Yes thanks...I was almost on the right track...thanks!

OpenStudy (zzr0ck3r):

0^0= 0^(1-1)=0^1/0^1=0/0=undefined